### v1

# Early release of low-frequency atlas of continuous gravitational waves

Abstract
We present the public release of the low-frequency atlas of continuous gravitational waves, covering signals with frequencies from 20 Hz to 200 Hz and frequency derivatives from $-5\times 10^{-11}$ to $5\times 10^{-11}$ Hz/s. Compared to the previous atlas releases, this version demonstrates significant improvements in sensitivity and sky resolution. In the most sensitive region even the worst-case upper limits on gravitational wave strain are below $10^{-25}$. The atlas data is being released ahead of the completion of the full follow-up analysis.
I Introduction
The first continuous gravitational wave atlas was released several years ago [2] and it covered signals with frequencies between 500-1000 Hz and frequency derivatives from $-5\times 10^{-11}$ to $5\times 10^{-11}$ Hz/s. Since then, we produced new atlases that progressively expanded parameter space coverage [2, 3, 4].
With this paper, we make the low-frequency atlas publicly available to improve coverage in the low frequency regime. The sky resolution has been enhanced by a factor of $6$ compared to [4], and the use of the full O3 data [5, 6] combined with a longer coherence length has improved the sensitivity of the search. In fact, the worst-case all-sky upper limits have decreased by 30%, and for the first time are below $10^{-25}$ at the most sensitive frequencies (Figure 1).
Our atlas provides upper limits on continuous gravitational waves based on measurements of power in LIGO [7] data. Additionally, we include signal-to-noise ratio (SNR) peaks and their associated parameters.
Figure 1: Gravitational wave intrinsic amplitude $h_{0}$ upper limits at 95% confidence as a function of signal frequency. The upper limits reflect the sensitivity of the search. We also plot latest LIGO/Virgo and Einstein@Home all-sky population average results [17, 16, 3] and Falcon population average proxy.
II Signal model
Continuous gravitational waves are expected from rapidly rotating neutron stars with sustained non-axisymmetric deformation $\varepsilon$. The intrinsic amplitude of the gravitational wave signal at a distance $d$ from the source is proportional to the deformation, increases with the square of the signal frequency $f$ and decreases with the inverse of the distance to the source $d$:
$h_{0}=\frac{4\pi^{2}GI_{zz}f^{2}\varepsilon}{c^{4}d},$ (1)
where $I_{zz}=10^{38}$ kg m2 is the moment of inertia of the star with respect to the principal axis aligned with the rotation axis. Gravitational waves carry away energy, gradually slowing down the pulsar’s rotation. Most known radio pulsars have frequency derivatives below with ±$\pm 5\times 10^{-11}$ Hz/s.
To compute the atlas data we assume a source which emits gravitational waves with the following frequency evolution in the Solar System baricenter
$f(t)=f_{0}+(t-t_{0})f_{1}+(t-t_{0})^{2}f_{2}/2.$ (2)
Here source frequency $f_{0}$ covers frequencies from $20$ to $200$ Hz, source frequency derivative is constrained by $|f_{1}|\leq 5\times 10^{-11}$ Hz/s and source second frequency derivative is within $|f_{2}|\leq 10^{-20}$ Hz/s2. We chose epoch $t_{0}=1253800000$ s  GPS (2019 Sep 29 13:46:22 UTC), which is in the middle of our data set.
Higher order frequency derivatives are not explicitly searched over, instead relying on intrinsic robustness of a search with coherence lengths of 2-days and using one year of data.
The signal model of equation 2 and the parameter space choice were optimized for isolated rotating neutron stars, but apply more generally, in particular neutron stars in wide-period binary orbits [8] and boson condensate sources [9, 10, 11, 12].
III Atlas data
The atlas [13] partitions the parameter space by frequency and sky position. The frequency space is divided into 50 mHz frequency bands. For each frequency band, we construct a grid of sky locations. The sky points are sufficiently dense that the upper limit values associated to the point hold for all sky positions in the Voronoi domain associated with the point. The Voronoi domain is constructed using the angular distance on the celestial sphere.
We set different types of upper limits, established based on different assumptions on the polarization of incoming signals. The worst-case upper limit is computed as maximum over all polarizations and is dominated by linearly polarized signals. Circularly polarized signals yield the most stringent upper limits (lower curve in Figure 1), typically more than a factor of $2$ lower than worst-case upper limits, because circularly polarized signals couple best with the detector. Upper limits for signals of arbitrary polarizations can be estimated using the sky-position-dependent coefficients provided in the atlas[14].
All upper limits are derived using the universal statistics algorithm [15] and are strict frequentist 95% confidence level upper limits. The 95% is actually a design target, and actual values are expected to have higher confidence depending on the underlying noise distribution.
The SNR values included in the atlas data represent peak values over the Voronoi domain associated with the given sky point and over the given frequency band. We also provide the value of the frequency and the polarization of the highest SNR result.
With this atlas we also introduce an estimate of the signal strength necessary for our search to “lock-on” to a signal from a given sky-position, i.e. to detect the signal and correctly estimate its frequency and polarization. This is the “location-specific polarization-average lock-level proxy”.
# $f$ $\dot{f}$ SNR UL/$h_{0}$ $h_{0}$/LL $\Delta f$ In
Hz Hz/s (atlas) % % mHz
3 $108.85716$1.46e-17 34.1 126.7 189.4 0.031 Yes
5 $52.80832$4.03e-18 259.9 135.2 364.6 -0.005 Yes
6 $145.33976$6.73e-09 8.2 19.8 638.0 -30.505 No
8 $189.96498$8.65e-09 8.3 58.8 215.0 0.832 No
10 $26.33144$8.5e-11 51.2 86.4 120.5 -23.734 No
11 $31.42470$5.07e-13 41.7 245.4 87.4 -10.257 Yes
12 $37.70750$6.25e-09 9.7 106.4 133.3 47.890 No
Table 1: This table shows the frequency parameters of the hardware-injected continuous wave signals and how they appear in the atlas data. We show all the hardware injections within 20-200 Hz range, including those outside of our search space, as indicated by the “In” column. We show our upper limit UL at the injection sky position and polarization parameter values, and how it compares with the injection amplitude $h_{0}$: if the upper limits are correct, UL/$h_{0}$ has to be $>100$%. $h_{0}$/LL shows the ratio of signal strain to the lock level, and we expect to detect signals when this ratio is $>100\%$. This is the case for the first two hardware-injected signals in range, but not for the third, as it can be seen from the large mismatch difference $\Delta f$ between the recovered frequency and the signal frequency. For detectable signals in our parameter space we expect the $\Delta f$ to be within $0.5$ mHz. We use the reference time (GPS epoch) $t_{0}=1253800000$ s (2019 Sep 29 13:46:22 UTC). The script spatial_index_example2.R provided with the atlas shows how to obtain this data from the atlas.
IV Population average proxies
The location-specific polarization-average lock-level proxy is defined as the signal strain level at which the frequency and polarization parameters of the SNR peak in the correspoding Voronoi domain are correctly reported in the atlas for 95% of signals. The primary intended use for the proxy is to assess when a limited parameter-space follow-up of SNR-peak parameters has merit, or when the entire Voronoi domain needs to be reanalyzed with a more sensitive search in the given frequency band.
Our location-specific polarization-average lock-level proxies are computed for each atlas entry from an estimate of the background noise and are insensitive to the presence of plausible strength signals. In our injection study 99% of injections changed the lock level by less than 1%. We validate these proxies using test signals distributed uniformly on the sky and with frequencies in the 100-200 Hz band. For each test signal, we identify the nearest point in the atlas and compare the SNR peak parameters for that point to those of the test signal. Figure 2 shows the percentage of signals with SNR peak frequencies within $0.5$ mHz of the actual value, as well as the percentage of signals with both frequency and polarization correctly recovered, as a function of signal strength. We verify that when the signal strength is at the lock-level, the recovered fraction of signals across a population of signals with arbitrary polarization is 95%. Figure 2 also shows that frequencies are recovered most accurately, while polarization recovery is less precise. This is because polarization reconstruction relies heavily on amplitude modulation, introducing significant degeneracy, especially for nearly circularly polarized signals (Figures 3 and 4).
Location-specific lock-level proxies can be used to estimate the sensitivity of our search for a population of signals coming from any direction in the sky, and this can in turn be used to generate an estimate of the so-called population-average upper limits over the whole sky, in the absence of signals. This is useful because most all-sky searches [17, 16] report as population-average upper limits. These are conceptually simple to implement by simulating search results from populations of signals and comparing with the observation results, until the signal amplitude that yields the 95% quantile is identified. However, practical challenges may arise, and the computational cost of these estimates is generally not trivial. The Falcon searches, like the one presented here, and its predecessor the Powerflux search, have used a different approach, establishing upper limits for fixed polarizations using the universtal statistics [15], which do not suffer from the same limitations. However, this can make comparison with the results of other pipelines more difficult.
With the lock-level proxies in hand, we provide an estimate of the population-average all-sky upper limits in the absence of signals, by taking the 95% quantiles of the location-specific lock-level proxies, resulting in a 90% confidence level. The resulting curve is shown in Figure 1, and it can be used for pipeline sensitivity comparisons. Its physical meaning should be interpreted as a level of strain at which 90% of signals are expected to have correct SNR peak parameters in the atlas, and not as actual upper limit results. Computationally expensive test-signal recovery simulations would be required to determine the actual confidence associated with the values in each band.
Figure 2: Signal recovery versus the ratio of signal strain to lock-level estimated on data with the test signals. The red line marks the $95$% threshold. The blue crosses show the percentage of points where the SNR peak frequency was withing $0.5$ mHz of the injected value.Given the $50$ mHz frequency band there is a 1% chance of false coincidence. The magenta circles show the percentage of points where $\iota$ was within $0.7/\sin(\iota)$ of injected value and $\psi$ was also within $0.7$ of injected value modulo $\pi/2$. Such wide tolerances are due to shoulders in $\iota$ and $\psi$ recovery (see Figures 3 and 4).
The all-sky population average upper limit proxy shown in Figure 1 are the 95% quantile of the location-specific population-average upper limit proxies, resulting in a 90% confidence level for the curve shown in Figure 1 which is meant to be used for pipeline comparison purposes only.
Figure 3: Recovery of the inclination angle $\iota$. We plot only points corresponding to signals with strains at or above the population average lock-level. The higher uncertainty for nearly circular polarized signals results in wide shoulders in the histogram.
Figure 4: Recovery of the polarization angle $\psi$ from the SNR peaks. We plot only points corresponding to injected strain at or above the polarization-average lock-level. The left plot restricts data to nearly linearly polarized signals. The right plot includes all points, but multiplies the difference between injected and recovered $\psi$ values (modulo $\pi/2$) by $\sin(\iota)$. This way the mismatch in $\psi$ is deweighted as $\iota$ approaches circular polarizations 00 or $\pi$ (the “poles” of the polarization sphere).
V Conclusions
We have conducted a wide-band search of the full O3 data covering frequencies from 20 to 200 Hz. The results have been compiled into an atlas, which we release while the final analysis is ongoing. We introduce population average lock-level proxies for every sky location in the atlas data. The atlas includes several examples, written in R, that can be a starting point for future searches.